Fill in the blanks below with decimal numbers (not fractions):

Suppose the demand for a product can be modeled with the equation P = 100 − 2Q and the corresponding supply curve can be modeled with P = 3Q.

In equilibrium, _____________units of the good will be sold at a price of $ ________each.

At $66, the quantity demanded is __________units.

At $51, the quantity supplied is ________units

The deadweight loss of a $66 price floor would be $_________
Answer #1

In equilibrium, quantity demanded is equal to the quantity supplied so that we have

100-2Q=3Q

Q=100/5= 20 units. Price= 3*20 = $60

Therefore in equilibrium 20 units of the good will be sold at a price of $60.

When the price is $66 quantity demanded is 17 units

When the price is $51 the quantity supplied is 17 units.

At a price floor of 66 dollar, quantity demanded is 17 units and quantity supplied is 22 units.

Deadweight loss is given by 0.5*(66-51)*(20-17) = 22.50